OpenStudy (faman39):

The volume of a cone is 300 cubic kilometers. If the radius of the base is 6 kilometers, what is the height of the cone? Use 3.14 for pi . Show your work and round to the nearest tenth.

OpenStudy (faman39):

@JakeV8 can you able help me with this?

OpenStudy (anonymous):

do you have a formula for volume ofa cone?

OpenStudy (faman39):

V=1/3pir^2h

OpenStudy (anonymous):

thanks :) So V = 300 = (1/3) π r^2 h = (1/3) π 6^2 h = (1/3) π (36)h And you are supposed to use 3.14 for π so.... 300 = (1/3)(3.14)(36)h Just solve for h

OpenStudy (faman39):

Thank you so much @JakeV8 :)

OpenStudy (faman39):

But i m little confuse still

OpenStudy (anonymous):

what were you having trouble with? Glad to help... this just seemed like straightforward substituting once you had the formula... but I didn't have the formula, you did :)

OpenStudy (faman39):

Dont know, why?

OpenStudy (faman39):

I think i didnt get the process.

OpenStudy (anonymous):

ok, no worries :) glad to help. You just start with the formula and realize that you have values for all the variables except height.

OpenStudy (faman39):

ok i think i got it bit now

OpenStudy (faman39):

Thanks alot @JakeV8 , i will find a way out, u helped alot always :) your truly a great !